OpenStudy (anonymous):

Part 1: Simplify the complex fractions.

OpenStudy (anonymous):



OpenStudy (anonymous):

Part 2: Are the complex fractions equivalent? Explain why or why not.

OpenStudy (anonymous):

they are. \[\frac{ a }{ b } \div \frac{ c }{ d } = \frac{ a }{ b } \times \frac{ d }{ c }\]

OpenStudy (anonymous):

I dont understand?

OpenStudy (anonymous):

Is that the answer?

OpenStudy (anonymous):

maybe not the answer your teacher wants. but it is a rule of how you divide 2 fractions. you multiply it by the reciprocal of the divider

OpenStudy (anonymous):